OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

y = –x – 1 y = –x + 5 y = x + 1 y = x + 5

OpenStudy (anonymous):

It is either the 3rd or 4th choice . Why?

OpenStudy (anonymous):

y=x+1 goes through (2,3)

OpenStudy (anonymous):